OpenStudy (anonymous):

multiply (x + 7)² x^2 + 49?

OpenStudy (anonymous):

am I correct?

OpenStudy (imtiaz7):

by using formula: (a+b)^2= a^2+2ab+b^2 so it will be: x^2+14x+49 here a=x and b=7

OpenStudy (anonymous):

are you sure?

OpenStudy (imtiaz7):

yes!

OpenStudy (anonymous):

(x+7)^2=x^2+14x+49

OpenStudy (triciaal):

@dontre no, |you forgot the middle term dw:1424843894990:dw|

OpenStudy (triciaal):